OpenStudy (anonymous):

Which expressions represent the distance between the two points? Choose all answers that are correct. The number line shows two points with coordinates negative 11 and positive 7. A. |–11 – 7| B. |–11 + 7| C. |7 – (–11)| D. –11 – 7

OpenStudy (anonymous):

@JoeDeWise

OpenStudy (anonymous):

On the number line, the distance between negative 11 and positive 7 is 18 units. Choice A: The absolute value of (-11 - 7) = absolute value of (-18) = 18. Choice C: The absolute value of [ 7 - (-11) ] = absolute value of (7+11) = 18. A and C are both the same number, and it's the correct answer to the question.

OpenStudy (anonymous):

thx so a and c

OpenStudy (anonymous):

Yes.